OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. –2x – y = –7 2x – y = –7 2x – y = 11 –2x + y = –11

OpenStudy (anonymous):

Answer is B. 2x - y = -7 or in standard form its y = 2x + 7

OpenStudy (anonymous):

Thank you , @mayayamfundo Can you help me with another one? Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form. y = 3x – 6 y = 3x – 3 y = 3x + 3 y = 3x + 6

OpenStudy (anonymous):

Now you need to find gradient or slope first here. I hope you know how to do this. So slope here is 3...

OpenStudy (anonymous):

So correct answer is B

OpenStudy (anonymous):

This is the hardest subject to me and it is all going in one ear coming out the other :/ Thank you so much.

OpenStudy (anonymous):

And for me this is the easiest stuff... Been there done that

OpenStudy (anonymous):

But you will be fine if you ask for help

OpenStudy (anonymous):

I have the test done, but I keep missing these three and having to restart it, do you mind helping me with one more? @mayayamfundo If not it is fine you have saved my life enough with just two questions :p Write the equation of the line that passes through (2, 2) and (6, 3) in standard form. 4x + 4y = 6 x + 4y = 12 4x + 4y = –12 x – 4y = –6

OpenStudy (anonymous):

Don't stress, I can do Maths all day non-stop, so I don't mind. Again here you have to find the slope using the two coordinates. So our slope in this case is 1/2

OpenStudy (anonymous):

So what do you do with the slope? Where do you plug it in at?

OpenStudy (anonymous):

You answer here is D

OpenStudy (anonymous):

This is kind of a tough one, but the easiest way is to plug in the given values for intercepts into each equation and see if they are true and only onw will be true. If you know what I mean

OpenStudy (anonymous):

For example. If we take the coordinates (2;2)... When you substitute x=2 and y=2 into equation D you get -6. Again when you plug in the second set of coordinates x=6 and y=3 you get -6. So that is how you check

OpenStudy (anonymous):

Kind of tough? lol I still am completely lost :o but I did work it out on paper the way you said so it gave me a little bit of a guide for the next one that I have to do, thank you so much